OpenStudy (anonymous):

A driver travelling from Waterford to Donegal wished to time his journey so that he can meet a friend for lunch at 1:15pm in Mullingar. here is the information to help you: The distance to Donegal is approximately. 360km. When he reaches Mullingar he will have completed 2/3 of his journey. He expects to travel at an average speed of 90kph. At what time should he leave? Please tell me how to do it:)

OpenStudy (anonymous):

Help!!!!!!!!!!!!!!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[r t = d, r = \frac{d}{t} \]Using Military time, ie: 1PM -> 1300\[\left(13+\frac{15}{60}\right)-\frac{\frac{2}{3}*360}{90}=\frac{127}{12}=10.5833\to 10:35\text{ AM} \]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Sorry, the solution was based on t = d/r, not r = d/t as I posted above. Thank you for the medal.